### v1

# FM4Com: Foundation Model for Scene-Adaptive Communication Strategy Optimization

Abstract
The emergence of sixth-generation (6G) networks heralds an intelligent communication ecosystem driven by AI-native air interfaces. However, current physical-layer designs—typically following modular and isolated optimization paradigms—fail to achieve global end-to-end optimality due to neglected inter-module dependencies. Although large language models (LLMs) have recently been applied to communication tasks such as beam prediction and resource allocation, existing studies remain limited to single-task or single-modality scenarios and lack the ability to jointly reason over communication states and user intents for personalized strategy adaptation. To address these limitations, this paper proposes a novel multimodal communication decision-making model based on reinforcement learning. The proposed model semantically aligns channel state information (CSI) and textual user instructions, enabling comprehensive understanding of both physical-layer conditions and communication intents. It then generates physically realizable, user-customized link construction strategies that dynamically adapt to changing environments and preference tendencies. A two-stage reinforcement learning framework is employed: the first stage expands the experience pool via heuristic exploration and behavior cloning to obtain a near-optimal initialization, while the second stage fine-tunes the model through multi-objective reinforcement learning considering bit error rate, throughput, and complexity. Experimental results demonstrate that the proposed model significantly outperforms conventional planning-based algorithms under challenging channel conditions, achieving robust, efficient, and personalized 6G link construction.
Index Terms:
Reinforcement learning, large language models, physical layer, modality alignment.
I Introduction
The forthcoming sixth-generation (6G) wireless networks herald a paradigm shift that transcends conventional connectivity. Central to this vision is the concept of AI-native air interfaces, which enable self-optimizing networks capable of understanding user intent and dynamically adapting to complex and extreme environments[1, 2, 3]. However, current mainstream physical-layer architectures still rely on modular, layer-wise optimization frameworks, making it challenging to achieve the desired level of intelligence and scene-adaptivity[4]. Conventional designs typically decompose the system into separate, scene-irrelevant functional modules optimized in isolation[4]: advanced channel coding schemes such as LDPC and Polar codes are tailored to specific channel models[5]; adaptive modulation and coding (AMC) algorithms adjust modulation orders based on channel state information (CSI)[6]; and MIMO precoding and beamforming techniques aim to maximize spectral efficiency[7]. Although these methods enhance individual module performance, their isolated optimization neglects the complex, nonlinear interdependencies among components, resulting in globally suboptimal performance and diminished end-to-end efficiency.
With their powerful contextual understanding, cross-modal semantic fusion, and global reasoning capabilities, large language models (LLMs) can simultaneously process heterogeneous information and generate globally consistent decisions, offering new possibilities for overcoming the fragmented optimization inherent in traditional communication system design[8]. Existing research applying LLMs to wireless communications can generally be categorized into two main directions.
The first line of research focuses on leveraging pretrained LLMs to enhance specific communication tasks. For example, [9] proposes an LLM-based downlink channel prediction method, designing task-specific embedding layers for the frequency and angular domains while keeping the LLM backbone frozen. In [10], an LLM-based beam prediction approach is proposed that utilizes textual prompts and achieves superior robustness and generalization compared with conventional LSTM-based models. In [11], LLMs are employed for resource allocation and, after fine-tuning on small-scale datasets, achieve performance comparable to advanced reinforcement learning algorithms. Our previous work [12] proposes a modality-aligned LLM for channel prediction, which effectively narrows the modality gap between CSI and linguistic knowledge to enhance prediction accuracy.
The second research direction focuses on developing task-specific foundation models tailored for wireless communication. For example, [13] proposes a task-independent universal channel embedding base model for communication by predicting the content of masked channel patches during training. Similarly, [14] considers the three dimensions of time, space, and frequency, and obtains a multi-task base model covering communication perception through pre-training and multi-task adaptive fine-tuning. [15] proposes a unified self-supervised framework that combines contrastive learning and masked reconstruction, specifically designed for multi-task channel representation learning. Despite these advances, notable limitations remain. Most existing efforts still focus on single tasks or modalities, leaving the core strengths of foundation models—deep multimodal reasoning, cross-modal fusion, and unified sequential decision-making—underexplored. This isolated usage pattern prevents communication systems from achieving unified situational awareness and globally optimal end-to-end decisions.
To address the aforementioned challenges, we propose a foundation model designed for interactive communication strategy customization with a Chain-of-Thought-enhanced Reinforcement Learning (CoT-RL) framework named IntraCom-FM, which enables the construction of end-to-end communication links that adapt to specific user preferences under varying channel conditions. The main contributions of this paper can be summarized as follows:
  • We propose a foundation model for scenario-adaptive communication strategy optimization, termed FM4Com. The model possesses the capability to comprehensively perceive both CSI and user intents. By semantically integrating physical-layer conditions with natural language instructions, FM4Com can generate customized and optimal link configuration strategies that align with the current channel dynamics and user preferences, thereby achieving personalized and adaptive optimization of communication systems.
  • We propose a chain-of-thought-enhanced two-stage reinforcement learning framework. In the first stage, a heuristic exploration algorithm is employed to expand the reinforcement learning experience pool, followed by behavior cloning via supervised learning to obtain a near-optimal initialization. In the second stage, the model is fine-tuned through reinforcement learning to further improve decision quality. Throughout the training process, the model also generates user preference representations as intermediate reasoning steps, forming an internal chain of thought that guides subsequent decision-making.
  • Experimental results demonstrate that the proposed FM4Com model can adaptively construct robust and efficient link strategies tailored to diverse user requirements under various challenging channel conditions, significantly outperforming existing planning-based algorithms.
II System model
II-A Full-link Communication Simulation System
We consider a 3GPP-compliant physical-layer communication link[16], as illustrated in Fig. 1. The system consists of several functional modules, including channel coding, spreading, modulation, power control, resource allocation, channel estimation, equalization, and HARQ. Each module adopts a specific strategy denoted as $c_{i}$, and the combination of all selected strategies forms a complete physical-layer transmission chain $\bm{a}=\{a_{1},a_{2},\ldots,a_{n}\}$.
The end-to-end transmission process of the communication link can be abstractly modeled as:
$\hat{\bm{x}}=f(\bm{x};\bm{a},\bm{H}),$ (1)
where $\bm{x}$ denotes the transmitted data, $\hat{\bm{x}}$ represents the recovered data at the receiver, and $\bm{H}$ is the CSI.
Based on the relationship between transmitted and received data, the overall performance of the physical-layer system can be comprehensively evaluated in terms of BER, system rate, and complexity.
II-B State Space Modeling
The selection of strategy combinations across modules in the communication link can be modeled as a RL process. Unlike conventional multi-step RL problems, the considered scenario corresponds to a single-step decision process, where each transmission involves a one-shot optimization of all module strategies.
We define the decision process as $M=\{\bm{S},\bm{A},r\}$, where $\bm{S}$ denotes the state space, $\bm{A}$ denotes the action space, and $r:\bm{S}\times\bm{A}\rightarrow\mathbb{R}$ represents the reward function. The objective is therefore to learn an optimal policy $\pi:\bm{S}\rightarrow\bm{A}$ that maximizes the expected immediate reward:
$\max_{\pi\in\Pi}\;\mathbb{E}_{\bm{s}\sim S,\,\bm{a}\sim\pi(a|s)}\big[r(\bm{s},\bm{a})\big].$ (2)
where $\bm{a}\in\bm{A}$ denotes the action and $\bm{s}\in\bm{S}$ denotes the state. For generality, $\pi(\bm{a}|\bm{s})\in[0,1]$ denotes the probability that policy $\pi$ selects action $\bm{a}$ given state $\bm{s}$.
In our framework, the proposed model receives both textual information and CSI as inputs, and outputs the optimal strategy for each module in the physical-layer communication chain. Accordingly, the policy parameterized by $\theta$ can be expressed as:
$\pi_{\theta}:\mathcal{O}_{\text{text}}\times\mathcal{O}_{\text{CSI}}\rightarrow\bm{a},$ (3)
where $\mathcal{O}_{\text{text}}$ denotes the textual observation space, $\mathcal{O}_{\text{CSI}}$ represents the channel state information space, and $\bm{c}$ corresponds to the set of strategies for all physical-layer modules.
III Proposed Method
In this section, we introduce the structure of our proposed FM4Com and the CoT-RL training method.
III-A Multimodal Communication Decision Large Language Model
As illustrated in Fig. 1, the FM4Com consists of a text encoder, a CSI encoder, a cross-modal attention module for textual semantic compression, a pretrained LLM, and a generator responsible for final strategy generation. The generator comprises $N_{a}$ actor networks, each designed to produce the optimal configuration strategy for a specific physical-layer module.
Figure 1: Our proposed FM4Com.
The input CSI sequence $\bm{H}\in\mathbb{R}^{L\times F}$ is processed by a CSI encoder to extract CSI features $\bm{{\mathit{Y}}}_{\text{csi}}$, where the CSI encoder adopts the same architecture as the preprocessing module in our previous work[11], which can be expressed as:
$\bm{{\mathit{Y}}}_{\text{csi}}=\text{Encoder}_{\text{csi}}(\bm{H})),$ (4)
where $\bm{{\mathit{Y}}}_{\text{csi}}\in\mathbb{R}^{L\times d}$.
Meanwhile, the user request $\bm{T}$ is converted into text tokens through a tokenizer and embedded into the same semantic space as the LLM. The embedding process can be formulated as:
$\bm{T}_{\text{text}}=\text{Tokenizer}(\bm{T}),$ (5)
where $\bm{T}_{\text{text}}\in\mathbb{R}^{L_{1}\times d}$ represents the word embeddeing of the original text.
Directly feeding long text sequences into the LLM leads to excessive computational overhead. To address this issue, we introduce a connector module before the text tokens are input into the LLM, which performs semantic filtering on textual embeddings. As illustrated in Fig. 1, the connector consists of three main stages: linear projection, self attention , and cross attention. First, a fully connected layer is employed to reduce the sequence length of text embeddings:
$\bm{Z}_{\text{text}}=\text{Linear}(\bm{T_{\text{text}}}),$ (6)
where $\bm{Z}_{\text{text}}\in\mathbb{R}^{L_{2}\times d}$.
Next, a self-attention mechanism is applied to establish semantic dependencies and aggregate global textual information:
$\tilde{\bm{Z}}_{\text{text}}=\text{Self-Attention}(\tilde{\bm{Z}}_{\text{text}}),$ (7)
Finally, inspired by the Q-Former structure, a cross-modal attention mechanism is adopted, where the CSI features act as queries to filter the textual embeddings semantically. Let $\bm{Y}_{\text{csi}}\in\mathbb{R}^{L\times d}$ denote the CSI representation aligned with the LLM’s semantic space. The cross-modal attention process can be formulated as:
$\bm{{\mathit{Y}}}_{\text{text}}=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{c}}\right)V,$ (8)
where
$Q=\bm{{\mathit{Y}}}_{\text{csi}}W_{Q},\quad K=\tilde{\bm{Z}}_{\text{text}}W_{K},\quad V=\tilde{\bm{Z}}_{\text{text}}W_{V},$ (9)
where $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}\in\mathbb{R}^{d\times d}$ are attention parameters. This operation allows CSI-driven queries to focus on task-relevant semantic tokens, effectively compressing long textual inputs into concise and context-aware embeddings.
After obtaining the CSI-aligned text features, we concatenate them with the projected CSI embeddings and feed them into the pretrained LLM backbone:
$\bm{{\mathit{Y}}}_{\text{LLM}}=\text{LLM}\left(\text{Concat}[\bm{{\mathit{Y}}}_{\text{text}};\bm{{\mathit{Y}}}_{\text{csi}}]\right),$ (10)
where $\bm{{\mathit{Y}}}_{\text{LLM}}\in\mathbb{R}^{2L\times d}$ denotes the hidden representation from the final transformer layer, and $L_{c}$ is the length of the compressed text sequence. Within the reasoning process, the LLM generates an intermediate preference vector $\bm{{\mathit{p}}}$ that reflects the user’s optimization tendency, such as throughput-oriented or complexity-sensitive objectives.
Finally, the preference vector $\bm{{\mathit{p}}}$ is concatenated with the multimodal LLM output and passed through multiple actor networks to produce the final decision strategies for different physical-layer modules. The output of the $i$-th actor network can be expressed as
$a_{i}=g_{\text{actor},i}\!\left(\text{Concat}[\bm{{\mathit{Y}}}_{\text{LLM}};\bm{{\mathit{p}}}]\right),\quad i=1,2,\ldots,N_{a},$ (11)
where $N_{a}$ denotes the number of decision modules such as channel coding, modulation, power control, and resource allocation. The overall physical-layer configuration is thus represented as
$\bm{a}=\{a_{1},a_{2},\ldots,a_{n}\},$ (12)
which forms the final multimodal communication strategy determined jointly by the semantic intent and the CSI conditions.
III-B CoT-RL Training Method
Figure 2: RL method.
The proposed model is trained using a two-stage CoT-RL procedure, which integrates behavior cloning for initialization and reinforcement learning for fine-tuning. The entire training process aims to optimize the multimodal decision model such that it can reason over both textual and CSI features to generate adaptive and high-performance communication strategies.
In the first stage, we perform behavior cloning to pretrain the model using a set of high-quality samples collected through a greedy search algorithm. Specifically, during early exploration, the multimodal model interacts with the environment to generate decision samples $(\bm{{\mathit{s}}},\bm{{\mathit{a}}},r)$, where $\bm{{\mathit{s}}}$ denotes the state composed of textual and CSI embeddings, $\bm{{\mathit{a}}}$ represents the selected strategy $\bm{{\mathit{a}}}$, and $r$ is the obtained reward. The greedy algorithm selects actions that yield the highest immediate reward, and the resulting tuples are stored in a replay buffer $\mathcal{D}$. The model is then trained to imitate these expert-like decisions through supervised learning:
$\mathcal{L}_{\text{BC}}=-\mathbb{E}_{(\bm{{\mathit{s}}},\bm{{\mathit{a}}})\sim\mathcal{D}}\big[\log\pi_{\theta}(\bm{{\mathit{a}}}|\bm{{\mathit{s}}})\big],$ (13)
where $\pi_{\theta}$ denotes the policy parameterized by the LLM-based actor and $\theta$ represents the trainable parameters. This behavior cloning stage enables the model to quickly learn a near-optimal policy by mimicking the expert trajectories in the buffer, effectively providing a warm start for subsequent reinforcement learning.
After behavior cloning, the model undergoes reinforcement learning fine-tuning to further optimize decision performance under the true environment dynamics. Given the state $\bm{{\mathit{s}}}$, the model samples an action $\bm{{\mathit{a}}}\sim\pi_{\theta}(\bm{{\mathit{a}}}|\bm{{\mathit{s}}})$, receives a scalar reward $r=f_{\text{reward}}(\text{BER},\text{Rate},\text{Complexity})$, and updates its parameters to maximize the expected return:
$J(\theta)=\mathbb{E}_{\bm{{\mathit{s}}},\bm{{\mathit{a}}}\sim\pi_{\theta}}[r(\bm{{\mathit{s}}},\bm{{\mathit{a}}})].$ (14)
The gradient of the objective can be estimated using the policy gradient method as
$\nabla_{\theta}J(\theta)=\mathbb{E}\big[\nabla_{\theta}\log\pi_{\theta}(\bm{{\mathit{a}}}|\bm{{\mathit{s}}})\cdot r(\bm{{\mathit{s}}},\bm{{\mathit{a}}})\big].$ (15)
To stabilize training, the behavior cloning loss is combined with the reinforcement learning objective to form the final joint loss:
$\mathcal{L}_{\text{total}}=-J(\theta)+\lambda\mathcal{L}_{\text{BC}},$ (16)
where $\lambda$ is a balancing coefficient controlling the influence of imitation learning on policy optimization.
The reward function is designed as a weighted linear combination of three key communication performance indicators: BER, Rate, and complexity. The overall reward is defined as
$r=w_{\text{ber}}\cdot R_{\text{ber}}+w_{\text{rate}}\cdot R_{\text{rate}}+w_{\text{comp}}\cdot R_{\text{comp}},$ (17)
where $w_{\text{ber}},w_{\text{rate}},w_{\text{comp}}\in[0,1]$ are weighting coefficients satisfying $w_{\text{ber}}+w_{\text{rate}}+w_{\text{comp}}=1$, and the normalized terms $R_{\text{ber}}$, $R_{\text{rate}}$, and $R_{\text{comp}}$ represent the contributions of BER, rate, and complexity, respectively.
To reflect user-specific communication preferences, these weights are adaptively determined based on the preference vector $\bm{{\mathit{p}}}$ inferred by the LLM:
$w_{\text{ber}},w_{\text{rate}},w_{\text{comp}}]=f_{\text{pref}}(\bm{{\mathit{p}}}),$ (18)
where users emphasizing reliability are assigned larger $w_{\text{ber}}$, while those focusing on high throughput or low complexity receive higher $w_{\text{rate}}$ or $w_{\text{comp}}$, respectively. Moreover, during the entire training process, we incorporate a cross-entropy loss $\mathcal{L}_{\text{pref}}$ on user preference prediction into the overall objective to ensure that the model accurately understands the user intent.
IV Numerical Results
This section presents our simulation results. First, we describe the simulation setup in detail. Then, we compare FM4Com with several existing policy selection methods to evaluate the effectiveness of the proposed approach. In addition, we present a human-computer interactive question-answering example to demonstrate the intelligence and flexibility of FM4Com.
IV-A Experimental Setup
We employ the QuaDRiGa channel generator to create a time-varying CSI dataset compliant with the 3GPP standard, covering multiple scenarios such as urban, rural, and highway environments. Corresponding textual descriptions are generated based on the characteristics of each scenario. To represent user preferences, we randomly assign three types of preference settings across all scenarios: low bit error rate, high throughput, and conventional systems. The GPT-2 model is adopted as the pretrained LLM backbone for validation.
IV-B Model Decision Performance
We first analyze the FM4Com’s ability to generate personalized communication strategies under different user preferences. For comparison, we employ three baseline methods for strategy selection: random strategy selection, greedy search, and beam search, with the beam width of the beam search method set to 3. As illustrated in Fig.3, the proposed model and baseline methods are evaluated across various channel environments in terms of BER, system complexity, and system rate.
(a) BER
(b) Rate
(c) Complexity
Figure 3: Performance comparison of different strategy selection methods for low BER strategy under variou SNRs.
For the low-BER preference, the FM4Com effectively captures the user’s intent and outputs robust communication strategies. As the SNR increases, the BER of all methods decreases; however, the reduction achieved by random selection is minimal, while the greedy and beam search algorithms demonstrate moderate improvements. In contrast, FM4Com significantly outperforms all baselines, achieving near-zero BER across a wide SNR range. This advantage arises because greedy and beam search methods only optimize local module-level performance. In contrast, FM4Com leverages reinforcement learning with end-to-end feedback, enabling globally optimal decisions across the entire communication chain. It is also observed that the remarkably low BER achieved by the proposed model comes at the cost of reduced throughput, which aligns with the user’s low-BER preference emphasizing reliability over efficiency.
(a) BER
(b) Rate
(c) Complexity
Figure 4: Performance comparison of different strategy selection methods for high rate strategy under various SNRs.
For the high-throughput preference, the FM4Com adapts its strategy accordingly, achieving higher throughput than in the low-BER case while maintaining a lower BER than random selection. Although greedy and beam search methods can yield higher throughput, their solutions are overly aggressive, leading to extremely high BER and rendering the constructed communication links nearly incapable of transmitting valid information. In contrast, our model achieves a better trade-off between throughput and reliability, balancing user preference with practical system stability.
Under the low-complexity preference, the FM4Com tends to adopt more conservative strategies, exhibiting similar behavior to the low-BER preference scenario. This is because low-BER configurations are inherently stable across different environments, and the model naturally gravitates toward such reliable solutions when computational simplicity is prioritized.
(a) BER
(b) Rate
(c) Complexity
Figure 5: Performance comparison of different strategy selection methods for general strategy under various SNRs.
IV-C Model Complexity
TABLE I: Inference time.
method inference time
Random 0.004s
FM4Com 0.46 s
Beam Search 36.36s
Greedy 13.55s
To assess the practical feasibility of deploying the proposed framework in real-world communication systems, we further analyze the algorithmic complexity of different strategy planning methods in a single planning task. As shown in Table I, heuristic-based algorithms such as the greedy and beam search methods exhibit significantly higher inference latency compared with the proposed FM4Com. This is primarily due to their sequential search and evaluation nature, where each potential action combination must be explicitly explored or locally optimized, resulting in exponential or at least polynomial growth in computational cost as the action space increases. In contrast, our FM4Com leverages end-to-end learning to implicitly capture the complex dependencies between modules, allowing it to generate globally optimized strategies through a single forward pass of the neural network. This substantial reduction in inference time demonstrates the scalability and practicality of the proposed model.
IV-D Model Interaction
Figure 6: FM4Com interaction examples.
In this section, we present a concrete human–machine interaction example to demonstrate the functional capabilities and intelligent characteristics of the proposed FM4Com. As illustrated in Fig.6, under a given communication condition, the user provides both the known environmental information and the specific transmission intent as multimodal inputs to the model.Upon receiving these inputs, FM4Com accurately interprets the user’s intent, comprehends the underlying communication context, and infers the user’s preference orientation. For example, according to the query shown in Fig.6, the model can infer that the user emphasizes reliability in the communication system. Based on this understanding, the model autonomously generates an optimized and interpretable communication strategy configuration, which aligns with both the current channel state and the user’s semantic preference. This example highlights the model’s strong semantic reasoning and adaptive decision-making capabilities, demonstrating its potential to serve as an intelligent agent for human–AI collaborative communication system design.
V Conclusion
In this paper, we proposed FM4Com, a foundation model for interactive communication strategy customization. FM4Com has been designed to comprehensively perceive both CSI and user intent. By integrating physical-layer conditions with natural-language instructions, the model first generates an intermediate chain of thought to infer user-specific preferences, and subsequently produces a customized and optimal link configuration strategy that aligns with both the current channel state and the inferred user intent, thereby enabling personalized and adaptive optimization of communication systems. During training, heuristic algorithms have been employed to collect high-quality experience samples for behavior cloning, followed by reinforcement learning fine-tuning to obtain the optimal policy. Extensive experimental evaluations have demonstrated that FM4Com significantly outperforms traditional planning-based algorithms under various challenging channel conditions, while flexibly generating optimal physical-layer strategies tailored to different user preferences and communication environments.